42.1 ml of air is at 24.1 °C. What is its volume at 77.1 °C?
1
Expert's answer
2018-03-01T08:45:45-0500
Charles' Law

V1/T1 = V2/T2

V2 = V1·T2/T1

V1 = 42.1 mL

T1 = 297.25 K

T2 = 350.25 K

V2 = 42.1mL·350.25K/297.25K = 49.6 mL